2,4,5,6,7,8,9,10,12,14


1
Expert's answer
2022-02-23T14:32:06-0500

Solution:

2,4,5,6,7,8,9,10,12,14

n=10

To find mean.

"\mu=\dfrac{\Sigma x_i}n=\dfrac{2+4+5+6+7+8+9+10+12+14}{10}
\\=\dfrac{77}{10}
\\=7.7"